# Confidence-Triggered Detection: Accelerating Real-time Tracking-by-detection Systems

### v3

@@ Abstract
Real-time object tracking necessitates a delicate balance between speed and accuracy, a challenge exacerbated by the computational demands of deep learning methods. In this paper, we propose Confidence-Triggered Detection (CTD), an innovative approach that strategically bypasses object detection for frames closely resembling intermediate states, leveraging tracker confidence scores. CTD not only enhances tracking speed but also preserves accuracy, surpassing existing tracking algorithms. Through extensive evaluation across various tracker confidence thresholds, we identify an optimal trade-off between tracking speed and accuracy, providing crucial insights for parameter fine-tuning and enhancing CTD’s practicality in real-world scenarios. Our experiments across diverse detection models underscore the robustness and versatility of the CTD framework, demonstrating its potential to enable real-time tracking in resource-constrained environments.
1 Introduction
Deep learning models are extensively utilized across diverse domains, including image analysis [14], natural language processing [17], and energy [26]. Image-related tasks encompass classification [6], segmentation [5], detection, image understanding [20]. Notably, object detection achieves remarkable performance on benchmarks such as ILSVRC [9, 22, 23], catalyzing the adoption of deep learning methodologies in video analysis, where videos are treated as sequential image sequences.
However, video analysis requires more than just analyzing individual frames. It’s crucial to establish relationships between frames [1, 19, 30, 31]. This is particularly important in the context of multiple object tracking (MOT), which has numerous applications in surveillance systems, intelligent robotics, and autonomous vehicles [13].
Most recent MOT approaches leverage a tracking-by-detection paradigm [2]. This involves three key steps: (1) object detection in each frame, (2) object tracking based on information from previous frames, and (3) data association using location and feature data from both detector and tracker [29]. While algorithms like SORT perform well in terms of precision and accuracy, they often suffer from ID switches and struggle with occlusions. DeepSORT addresses this with a better association metric, but real-world deployment necessitates both high accuracy and real-time performance [30].
@@ 2 Related Work
2.1 Tracking-by-detection
The dominant paradigm in MOT is tracking-by-detection [3]. This approach leverages two key components: an object detector and a tracker. In each frame, the detector identifies and localizes objects. The tracker then predicts the current location of previously detected objects based on their past states. Finally, an association step links detections from the current frame with corresponding tracks using information from both the detector and tracker [30]. This approach offers improved accuracy compared to alternatives, as the tracker’s predictions can be continuously refined by up-to-date detections.
2.2 Real-time Object Tracking
Defining a universal speed threshold for real-time performance in MOT is challenging due to variations in video frame rates captured by different cameras. However, a system is generally considered real-time if its processing speed surpasses the input video frame rate [16]. If the processing speed falls behind, a delay accumulates, causing the system output to become out of sync with the actual video and events.
Common strategies to achieve real-time learning include model optimization [34], hardware optimization [12, 35, 8], and new training and evaluation strategy [32]. The new training and evaluation strategy is a common way for tracking, such as frame skipping, where the system processes only a subset of video frames. One approach involves skipping a fixed number of frames after each tracking assignment. However, this method can lead to missing critical frames, especially for fast-moving objects, resulting in significant accuracy drops.
Our proposed CTD approach also incorporates frame skipping. However, unlike fixed-interval skipping, CTD leverages a confidence score to actively trigger new detections. This allows the system to strategically skip frames with minimal object movement while maintaining tracking accuracy. We continuously run the tracker in all frames to monitor the discrepancy between the tracker’s predictions and the latest detections. This difference serves as a measure of confidence, and a significant discrepancy triggers a new detection to correct potential biases before they accumulate significantly. This approach aims to achieve a balance between processing speed and tracking accuracy compared to traditional fixed-interval frame skipping.
2.3 Lightweight Detectors
@@ 2.3.1 YOLOv3 Tiny
2.3.2 MobileNet-SSD
MobileNets are neural networks that perform efficiently. It can also be used on mobile devices and reach a fairly high accuracy [11]. SSD [21] uses VGG16 [27] to extract feature maps. SSD classifies and locates objects in a single forward pass. MobileNet-SSD combines SSD and MobileNets which perform high speed and relatively high accuracy. Using MobileNet SSD, the system reaches 17.5% MOTA and performs 15.604 Hz of speed.
2.3.3 SqueezeNet
@@
2.4 Deep SORT
Deep SORT [30] is an extension of the SORT algorithm [1] that incorporates appearance information for object matching. This enables Deep SORT to track objects even during extended occlusions. Additionally, it utilizes the Mahalanobis distance metric [4] to incorporate motion information into the association process.
@@ 3 Method
3.1 Tracking-by-detection Framework
@@
Confidence-Triggered Detection (CTD): This innovative method integrates the Mahalanobis Distance and a derived Confidence Score with Detector Trigger Modules for object detection. The Confidence Score, deduced from the Mahalanobis Distance using a chi-squared distribution, dictates whether to maintain tracking or invoke the detector for new object detection, as elaborated in Section 3.3. Specifically, tracking continuity is preserved when the score surpasses a predefined threshold, leveraging the Hungarian Assignment for consistent object association across frames. Conversely, the score below the threshold or instances exceeding the maximum frame skip threshold prompt the activation of the detector to refresh the object identification process. The intricacies of the CTD approach, including detector operations and decision mechanisms, are further expounded in Section 3.2.
3.2 Confidence-Triggered Detection (CTD) Approach
@@
Confidence Score Evaluation: The Mahalanobis Distance is calculated between two bounding boxes. The predicted bounding of the current frame by the Kalman filter and bounding boxes of the last frame are detected by the detector. By applying a chi-squared distribution to the Mahalanobis Distance, as outlined in Section 3.3, we obtain a confidence score, which will be used to determine if triggering a new round of detection in the Detection Trigger Module.
@@
3.3 Confidence Score Inference
@@
where denotes the information at the -th frame and denotes the information at -th frame. The is the matrix of all the predicted bounding boxes at the -th frame. The and represent the mean and covariance of detected bounding boxes. The Mahalanobis distance follows a chi-square distribution [4]. Given probability level and degrees of freedom , we can calculate the distance threshold corresponding to that probability using the inverse cumulative distribution function (CDF) [24, 30]:
@@
4 Experiment
| Method | MOTA | MT | ML | FP | FN | Speed (Hz) |
| TraByDetNs | 22.2 | 43 | 345 | 4,664 | 41,676 | 8.1 |
@@ 4 Experiment
| TSDA_OAL | 18.6 | 68 | 305 | 16,350 | 32,853 | 19.7 |
| GMPHD | 18.5 | 28 | 399 | 7,864 | 41,766 | 19.8 |
| LDCT | 4.7 | 82 | 234 | 14,066 | 32,156 | 20.7 |
| Ours(CTD) | 19.5 | 37 | 389 | 5,115 | 43,059 | 21.4 |
@@
4.1 System Setup
We developed a tracking-by-detection system following the framework outlined in Section 3.1. DeepSORT was chosen as the tracker due to its real-time capabilities, occlusion handling, and competitive MOTA scores [30]. YOLOv3 Tiny [25] served as the object detector due to its balance between speed and accuracy. DeepSORT integrated with YOLOv3 Tiny achieved the best processing speeds compared to the existing algorithms.
@@ 4.1 System Setup
4.2 CTD in Action: Frame-by-Frame Analysis
@@
4.3 Comparison with Existing Methods
Following an exhaustive exploration of multiple confidence score thresholds, we determined that establishing the threshold at 30% achieves a commendable equilibrium between precision and computational efficiency. Subsequently, we conducted comparative analyses between our proposed method CTD and existing non-skippable and skippable tracking approaches. The results are summarized in Table 1. As a non-skippable approach, TraByDetNs operates object detection for each frame, contrasting with skippable methods such as TSDA_OAL [10], TC_SIAMESE [33], and GMPHD [28], which perform detection intermittently across frames. Our proposed method, CTD, outperforms all evaluated methods in terms of processing speed and achieves superior MOTA (Multi-Object Tracking Accuracy), and FP (False Positive) metrics compared to other skippable methods. Furthermore, CTD’s performance in terms of MT (Mostly Tracked targets), and ML (Mostly Lost targets) is comparable to the best-performing skippable methods, underscoring its efficacy in balancing tracking accuracy with computational efficiency.
4.4 Trade-Off Between Accuracy and Speed
@@
Figure 3(c) illustrates the direct relationship between speed (FPS) and accuracy (MOTA). It highlights that CTD consistently achieves superior accuracy levels compared to the fixed frame skipping strategy (0% confidence threshold), underscoring the efficacy of CTD. Furthermore, it offers valuable guidance for selecting an appropriate confidence threshold. By referring to Figure 3(c), one can determine the confidence threshold associated with the highest speed while maintaining the desired level of accuracy.
@@
4.5 CTD with Different Detectors
The CTD framework’s adaptability was assessed across various detection models, such as YOLOv3, MobileNet, and SqueezeNet. According to the data in Table 2, a consistent speed-accuracy balance is maintained across all models when incorporated with CTD. Notably, a mere 10% reduction in the confidence score threshold to 90% achieves a significant speed enhancement of about 30%, with only a negligible 3% compromise in accuracy. When the confidence score threshold is further relaxed to 10%, there is an impressive increase in speed by approximately 145%, with an acceptable decrease in accuracy of 27%. These outcomes highlight two principal advantages of the CTD framework: its plug-and-play compatibility with different platforms and models, and its ability to deliver substantial improvements in processing speed while sustaining satisfactory accuracy levels.
5 Conclusion

### v4

# Index Terms: Multiple objects tracking, Object detection, Real-time trackingI Introduction

@@ Abstract
Real-time object tracking necessitates a delicate balance between speed and accuracy, a challenge exacerbated by the computational demands of deep learning methods. In this paper, we introduce Confidence-Triggered Detection (CTD), a novel approach that strategically skips object detection for frames exhibiting high similarity, leveraging tracker confidence scores. CTD not only enhances tracking speed but also preserves accuracy, surpassing existing tracking algorithms. Through extensive evaluation across various tracker confidence thresholds, we identify an optimal trade-off between tracking speed and accuracy, providing crucial insights for parameter fine-tuning and enhancing CTD’s practicality in real-world scenarios. Furthermore, our experiments across diverse detection models underscore the robustness and versatility of the CTD framework, demonstrating its potential to enable real-time tracking in resource-constrained environments.
Index Terms:
Multiple objects tracking, Object detection, Real-time trackingI Introduction
The field of object detection has witnessed a dramatic shift in recent years, with deep learning architectures achieving unprecedented performance. However, video analysis presents a unique challenge that goes beyond the analysis of isolated frames. Unlike static images, videos possess an inherent temporal dimension, where the relationships between frames hold crucial information. Capturing these temporal dynamics is essential for a comprehensive understanding of the visual content in videos [1, 2, 3, 4, 5, 6, 7]. This is particularly important in the context of multiple object tracking (MOT), which has numerous applications in surveillance systems, intelligent robotics, and autonomous vehicles [8, 9, 10, 11].
Most recent MOT approaches leverage a tracking-by-detection paradigm [12, 13]. This involves three key steps: (1) object detection in each frame, (2) object tracking based on information from previous frames, and (3) data association using location and feature data from both detector and tracker [14]. While algorithms like SORT perform well in terms of precision and accuracy, they often suffer from ID switches and struggle with occlusions. DeepSORT addresses this with a better association metric, but real-world deployment necessitates both high accuracy and real-time performance [3].
@@ II Related Work
II-A Tracking-by-detection
The dominant paradigm in MOT is tracking-by-detection [15, 16, 17, 18, 19]. This approach leverages two key components: an object detector and a tracker. In each frame, the detector identifies and localizes objects. The tracker then predicts the current location of previously detected objects based on their past states. Finally, an association step links detected objects from the current frame with corresponding tracks using information from both the detector and tracker [3, 20, 21, 22, 23]. This approach offers improved accuracy compared to alternatives, as the tracker’s predictions can be continuously refined by up-to-date detection.
II-B Real-time Object Tracking
Defining a universal speed threshold for real-time performance in MOT is challenging due to variations in video frame rates captured by different cameras. However, a system is generally considered real-time if its processing speed surpasses the input video frame rate [24, 25, 26]. If the processing speed falls behind, a delay accumulates, causing the system output to become out of sync with the actual video and events.
Frame skipping is a common way for real-time object tracking systems where the system processes only a subset of video frames. One approach involves skipping a fixed number of frames after each tracking assignment. However, this method can lead to missing critical frames, especially for fast-moving objects, resulting in significant accuracy drops.
Our proposed CTD approach also incorporates frame skipping. However, unlike fixed-interval skipping, CTD leverages a confidence score to selectively trigger new detection. This allows the system to strategically skip frames with minimal object movement while maintaining tracking accuracy. We continuously run the tracker in all frames to monitor the discrepancy between the tracker’s predictions and the latest detection. This difference serves as a measure of confidence, and a significant discrepancy triggers a new detection to correct potential biases before they accumulate significantly. This approach aims to achieve a balance between processing speed and tracking accuracy compared to traditional fixed-interval frame skipping.
II-C Lightweight Detectors
@@ II-C1 YOLOv3 Tiny
II-C2 MobileNet-SSD
MobileNets are neural networks that perform efficiently. It can also be used on mobile devices and reach a fairly high accuracy [29]. SSD [30, 31, 32] uses VGG16 [33] to extract feature maps. SSD classifies and locates objects in a single forward pass. MobileNet-SSD combines SSD and MobileNets which perform high speed and relatively high accuracy. Using MobileNet SSD, the system reaches 17.5% MOTA and performs 15.604 Hz of speed.
II-C3 SqueezeNet
@@
II-D DeepSORT
DeepSORT [3] is an extension of the SORT algorithm [1] that incorporates appearance information for object matching. This enables Deep SORT to track objects even during extended occlusions. Additionally, it utilizes the Mahalanobis distance metric [35, 36] to incorporate motion information into the association process.
@@ III Method
III-A Tracking-by-detection Framework
@@
Confidence-Triggered Detection (CTD): This innovative method integrates the Mahalanobis Distance and a derived Confidence Score with Detector Trigger Modules for object detection. The Confidence Score, deduced from the Mahalanobis Distance using a chi-squared distribution, dictates whether to maintain tracking or invoke the detector for new object detection, as elaborated in Section III-C. Specifically, when the confidence score falls below a predefined threshold or the number of consecutively skipped frames exceeds a predetermined threshold, new object detection will be triggered and the previous detection result will be replaced. Otherwise, object detection is skipped for this frame. The intricacies of the CTD approach, including detector operations and decision mechanisms, are further expounded in Section III-B.
III-B Confidence-Triggered Detection (CTD) Approach
@@
Confidence Score Evaluation: The Mahalanobis Distance is calculated between two bounding boxes. The predicted bounding of the current frame by the Kalman filter and bounding boxes of the last frame are detected by the detector. By applying a chi-squared distribution to the Mahalanobis Distance, as outlined in Section III-C, we obtain a confidence score, which will be used to determine if triggering a new round of detection in the Detection Trigger Module.
@@
III-C Confidence Score Inference
@@
where denotes the information at the -th frame and denotes the information at -th frame. The is the matrix of all the predicted bounding boxes at the -th frame. The and represent the mean and covariance of detected bounding boxes. The Mahalanobis distance follows a chi-square distribution [35]. Given probability level and degrees of freedom , we can calculate the distance threshold corresponding to that probability using the inverse cumulative distribution function (CDF) [37, 3, 38, 39, 40, 41]:
@@
IV Experiment
| Method | MOTA | FP | Speed (Hz) |
| TraByDetNs | 22.2 | 4,664 | 8.1 |
@@ IV Experiment
| TSDA_OAL | 18.6 | 16,350 | 19.7 |
| GMPHD | 18.5 | 7,864 | 19.8 |
| LDCT | 4.7 | 14,066 | 20.7 |
| Ours(CTD) | 19.5 | 5,115 | 21.4 |
@@
IV-A System Setup
We developed a tracking-by-detection system following the framework outlined in Section III-A. DeepSORT was chosen as the tracker due to its real-time capabilities, occlusion handling, and competitive MOTA scores [3]. YOLOv3 Tiny [27] served as the object detector due to its balance between speed and accuracy. DeepSORT integrated with YOLOv3 Tiny achieved the best processing speeds compared to the existing algorithms.
@@ IV-A System Setup
IV-B CTD Tracking: Frame-by-Frame Analysis
@@
IV-C Comparison with Existing Methods
Following an exhaustive exploration of multiple confidence score thresholds, we determined that establishing the threshold at 30% achieves a commendable equilibrium between precision and computational efficiency. Subsequently, we conducted comparative analyses between our proposed method CTD and existing non-skippable and skippable tracking approaches. The results are summarized in Table I. As a non-skippable approach, TraByDetNs operates object detection for each frame, contrasting with skippable methods such as TSDA_OAL [43], TC_SIAMESE [44], and GMPHD [45], which perform detection intermittently across frames. Our proposed method, CTD, outperforms all evaluated methods in terms of processing speed and achieves superior MOTA (Multi-Object Tracking Accuracy), and FP (False Positive) metrics compared to other skippable methods.
IV-D Trade-Off Between Accuracy and Speed
@@
Figure 3(c) illustrates the direct relationship between speed (FPS) and accuracy (MOTA). It highlights that CTD consistently achieves superior accuracy levels at the same speed compared to the fixed frame skipping strategy (0% confidence threshold), underscoring the efficacy of CTD. Furthermore, it offers valuable guidance for selecting an appropriate confidence threshold. By referring to Figure 3(c), one can determine the confidence threshold associated with the highest speed while maintaining the desired level of accuracy.
@@
IV-E A Comparison between CTD with Different Trackers
The CTD framework’s adaptability was assessed across various detection models, such as YOLOv3, MobileNet, and SqueezeNet. According to the data in Table II, a consistent speed-accuracy balance is maintained across all models when incorporated with CTD. Notably, a mere 10% reduction in the confidence score threshold to 90% achieves a significant speed enhancement of about 30%, with only a negligible 3% compromise in accuracy. When the confidence score threshold is further relaxed to 10%, there is an impressive increase in speed by approximately 145%, with an acceptable decrease in accuracy of 27%. These outcomes highlight two principal advantages of the CTD framework: its plug-and-play compatibility with different platforms and models, and its ability to deliver substantial improvements in processing speed while sustaining satisfactory accuracy levels.
V Conclusion